how do you find a proportion. 12% of 17
Answer by stanbon(75887)

how do you find a proportion. 12% of 17
---
x/12% = 17/100%
----
x = 12%[17/100%)
---
x = (12/100)*17
---
x = 0.12*17
----
x = 2.04
=============
Cheers,
Stan H.
==============